OpenStudy (anonymous):

If an equation is an identity, then how many solutions does it have? zero one infinite

OpenStudy (anonymous):

1

OpenStudy (anonymous):

?

OpenStudy (zzr0ck3r):

tax(x) = cos(x)/sin(x) would have infinite solutions

OpenStudy (kira_yamato):

infinite, since it holds true for all variables

OpenStudy (kira_yamato):

E.g. sin^2 x = 1 - cos^2 x has infinite # solution

OpenStudy (zzr0ck3r):

tan(x) = sin(x)/cos(x) would have infinite solutions rather, is this what your asking?